an instructor measured quiz scores and the number of hours studying among a sample of 20 college students. if ssxy=42, ssx=91, my=7, and mx=4, then what is the regression equation for this sample?
Answer #1

SSxy = 45

SSx = 91

Mx = 7

My = 4

Regression equation is given as:

y = b0 + b1*x

Formula for b1 = SSxy/SSx

b1 = 45/91 = 0.4945

Formula for b0 = My - b1*Mx

b0 = 4 - 0.4945*7

b0 = 0.5385

Thus, regression equation can be written as:

y = 0.5385 + 0.4945*x